OpenStudy (hatahateo):

Help! I'll fan and medal Find a cubic function with the given zeros. Square root of two., - Square root of two., -2

OpenStudy (hatahateo):

@sammixboo

OpenStudy (hatahateo):

Answers: f(x) = x3 + 2x2 - 2x + 4 f(x) = x3 + 2x2 + 2x - 4 f(x) = x3 - 2x2 - 2x - 4 f(x) = x3 + 2x2 - 2x - 4

sammixboo (sammixboo):

Square root of 2, and square root of -2?

OpenStudy (hatahateo):

\[\sqrt{2} , -\sqrt{2} , and -2\]

TheSmartOne (thesmartone):

If a function has roots of a, b, and c (all with multiplicity 1. Haven't heard of multiplicity? then don't worry about it.) the equation of the cubic function would be: f(x) = (x - a)(x - b)(x - c)

TheSmartOne (thesmartone):

and then you multiply (x-a)(x-b)(x-c) to get your final answer.

OpenStudy (hatahateo):

I'm just confused about what I put where..

TheSmartOne (thesmartone):

If a is a root, then the equation is (x - a) is sqrt(2) is a root then the equation is (x - sqrt(2)) do the same for the other two roots, and then you expand them

OpenStudy (hatahateo):

@sammixboo I don't understand..

TheSmartOne (thesmartone):

what don't you understand?

OpenStudy (hatahateo):

Ugh..All of it, I know you just explained it all but I've literally been doing math all day my brain is fried.

TheSmartOne (thesmartone):